OpenStudy (anonymous):

1 over the cube root of the quantity of x to the negative sixth power

OpenStudy (anonymous):

Well first off, a negative power means to take the reciprocal. So 1x−m−−−−√n=1x−√n−m=x−√nm And the root becomes the den'r in the rational exponent: x−√nm=xmn

OpenStudy (anonymous):

I got it right!! thanks :)

OpenStudy (anonymous):

Im sorry about the question marks, lol !!!

OpenStudy (anonymous):

haha no problem thank you so much, have a great day!!

OpenStudy (anonymous):

Yeah no problem.

OpenStudy (librarian):

Metals for everybody (=